# Enhancing Robot Navigation Policies with Task-Specific Uncertainty Management

### v2

@@
Abstract.
Robots performing navigation tasks in complex environments face significant challenges due to uncertainty in state estimation arising from sensor noise, environmental changes, and incomplete information. Effectively managing this uncertainty is crucial, but the optimal approach varies depending on the specific details of the task: different tasks may require varying levels of precision in different regions of the environment. For instance, a robot navigating a crowded space might need precise localization near obstacles but can operate effectively with less precise state estimates in open areas. This varying need for certainty in different parts of the environment, depending on the task, calls for navigation policies that can adapt their uncertainty management strategies based on task-specific requirements.
In this paper, we present a framework for integrating task-specific uncertainty requirements directly into navigation policies. We introduce the concept of a Task-Specific Uncertainty Map (TSUM), which represents acceptable levels of state estimation uncertainty across different regions of the operating environment for a given task. Using TSUM, we propose Generalized Uncertainty Integration for Decision-Making and Execution (GUIDE), a policy conditioning framework that incorporates these uncertainty requirements into the robot’s decision-making process.
We find that conditioning policies on TSUMs not only provides an effective way to express task-specific uncertainty requirements but also enables the robot to reason about the context-dependent value of certainty and adapt its behavior accordingly. We show how integrating GUIDE into reinforcement learning frameworks allows the agent to learn navigation policies that effectively balance task completion and uncertainty management without the need for explicit reward engineering. We evaluate GUIDE at scale on a variety of real-world robotic navigation tasks and find that it demonstrates significant improvements in task completion rates compared to baseline methods that do not explicitly consider task-specific uncertainty. Evaluation videos can be found at https://guided-agents.github.io.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Consider an autonomous surface vehicle (ASV) tasked with mapping underwater topography near a busy harbor. Performing such a task in a marine environment is often complicated by factors like waves and environmental changes. These elements can introduce significant uncertainty in the ASV’s state estimation. To effectively carry out its mission, the ASV needs to manage this uncertainty. Importantly, constantly striving to reduce uncertainty is not always necessary for task completion. When navigating through areas with heavy boat traffic or obstacles, precise localization is crucial to avoid collisions and collect accurate data. However, while traversing open waters during this mission, the ASV can tolerate higher positional uncertainty, allowing it to focus on efficient task execution rather than achieving pinpoint accuracy. This scenario underscores a fundamental challenge in robotic navigation: the acceptable level of uncertainty at any location in the operation space is inherently task-specific and context-dependent.
@@ 1. Introduction
Figure 1. Illustration of our method applied to a robotic navigation task. Top Left: The ASV is assigned a navigation task. Top Right: GUIDE interprets the task and generates a representation highlighting areas where the ASV needs higher positional certainty (dark blue). Bottom: The policy executed by the ASV. The white line represents the trajectory taken by the ASV towards the dock (green rectangle), the red dots indicate locations where the ASV actively reduces its state estimation uncertainty to satisfy task-specific requirements.
Traditional approaches to uncertainty management in robotic navigation often fall into two categories: minimizing uncertainty universally (Patle et al., 2019; Gonzalez-Garcia and Castañeda, 2021; Mavrogiannis et al., 2023; Gul et al., 2019) or enforcing fixed uncertainty thresholds (Cook and Zhang, 2020; Hanheide et al., 2017; Rana et al., 2020). Such one-size-fits-all strategy may be effective in controlled settings, they lack the flexibility to adapt to tasks that demand varying degrees of uncertainty management. For instance, in the ASV’s mission introduced earlier, strictly enforcing precise localization may be unnecessary when navigating open waters, where higher uncertainty does not hinder task execution.
@@
These existing methods often neglect the varying uncertainty requirements dictated by different contexts and tasks. While some approaches attempt to balance uncertainty reduction with task performance (Taylor et al., 2021; Spaan et al., 2015), they typically apply this trade-off uniformly across all situations, failing to account for task-specific needs that may vary within a single mission. Furthermore, developing effective uncertainty management strategies often require extensive reward engineering to capture the intricacies of different tasks (Zhu et al., 2021; Chen et al., 2019).
The fundamental issue lies in the disconnect between task specifications and uncertainty management. Current frameworks often treat these elements separately, resulting in sub-optimal performance when effective task completion demands varying uncertainty levels across different regions in the operating environment.
To address this disconnect, we pose the following question: How can robots reason about the uncertainty needs of specific tasks to guide their navigation policies?
@@
Contributions: With the aim of bridging the gap between high-level task descriptions and low-level uncertainty handling, we propose the concept of a Task-Specific Uncertainty Map (TSUM). A TSUM represents the acceptable levels of state estimation uncertainty across different regions of the environment, given a specific task. By spatially encoding the varying importance of uncertainty reduction, TSUMs provide a task-oriented representation of where precision is most critical for successful task completion.
@@
Building upon TSUM, we introduce Generalized Uncertainty Integration for Decision-Making and Execution (GUIDE), a policy conditioning framework that incorporates TSUMs into navigation policies. The core proposition of GUIDE is that by incorporating the acceptable levels of state estimation uncertainty across different regions of the environment into the navigation policy, robots can better reason about the context-dependent value of certainty. This enables the generation of policies that adapt their behavior based on the specific uncertainty requirements of the given task.
@@
The GUIDE framework is designed to be simple, flexible, and broadly applicable across various robotic domains. We demonstrate its effectiveness in real-time, real-world applications, focusing on marine autonomy using ASVs. Our experimental results show significant improvements in task performance when compared to state-of-the-art methods.
2. Related Works
Uncertainty Modeling in Robotic Navigation:
@@ 2. Related Works
Uncertainty management is a fundamental aspect of robotic navigation, where robots must make decisions based on imperfect information about their state and the environment (Patle et al., 2019; Schwarting et al., 2018). Probabilistic techniques such as Bayesian filters (Gustafsson et al., 2002a; Ko and Fox, 2009; Murphy, 1999) and particle filters (Gustafsson et al., 2002b; Thrun, 2002; Fox et al., 2001) have been widely used to enable robust localization and mapping in stochastic environments. These methods allow robots to estimate their position and navigate effectively despite sensor noise and environmental uncertainties. However, traditional approaches often treat uncertainty uniformly across the environment, without considering how different regions may require varying levels of certainty depending on the navigation task.
Reinforcement Learning for Navigation:
@@
Reinforcement Learning (RL) has been successfully applied to robotic navigation tasks, enabling agents to learn navigation policies through interaction with the environment (Wang et al., 2024; Xu et al., 2023; Zhu, 2023). Standard RL algorithms have been used to train robots to navigate in various settings, from simulated environments to real-world scenarios (Lee and Yusuf, 2022; Zhou et al., 2022; Wang et al., 2022). While these methods can learn effective policies in controlled environments (Zieliński and Markowska-Kaczmar, 2021; Patel et al., 2021; Kulhánek et al., 2021), they often perform suboptimally in real-world scenarios due to factors like partial observability, dynamic changes, and varying levels of uncertainty (Curtis et al., 2024; Morad et al., 2023). Additionally, standard RL typically does not account for uncertainty explicitly, which can hinder performance in complex environments where uncertainty plays a significant role.
@@
Figure 2. Overview of GUIDE. The framework consists of two phases: Embedding Alignment (Pretraining) and Deployment. In the embedding alignment phase, semantic and spatial embeddings are learned and aligned using triplet loss and attention mechanisms. During deployment, TSUMs are generated from natural language task descriptions and environmental data, and navigation policies are conditioned on these TSUMs to manage uncertainty in a task-aware manner.
Task-Specific Navigation Policies:
@@
Recent advances have focused on developing navigation policies that are conditioned on specific tasks or goals (Rakelly et al., 2019; Forestier et al., 2022; Nair et al., 2018a; Lynch et al., 2020). Task-conditioned RL allows agents to adapt their behavior based on high-level task specifications, improving flexibility and generalization. In navigation, this has been explored by training policies capable of handling objectives, adjusting the robot’s behavior according to the task (Zhu et al., 2017; Wang et al., 2019; Nair et al., 2018b). These approaches often neglect the varying importance of uncertainty management (Florensa et al., 2018; Kretzschmar et al., 2016) across different tasks and environments.
Uncertainty-Aware Reinforcement Learning in Navigation:
@@
Incorporating uncertainty into RL for navigation has been investigated to enhance exploration, robustness, and safety (Singi et al., 2024; Zhu, 2023). Some approaches introduce uncertainty penalization directly into the reward function, encouraging agents to avoid high-uncertainty regions (Guo et al., 2021; Zhu and Zhang, 2021; Li et al., 2024). Others utilize Bayesian RL methods to model uncertainty in value estimation, improving decision-making under uncertainty (Budd et al., 2023; Alali and Imani, 2024). Bootstrapped ensembles maintain multiple value function estimates to capture epistemic uncertainty, leading to more informed exploration strategies (Osband et al., 2016; Bai et al., 2022; Kumar et al., 2019a; Yarats et al., 2021). While these methods consider uncertainty, they often do so globally or uniformly (Paulig and Okhrin, 2024; Lan et al., 2023; Ovalle and Lucas, 2020) across the environment and do not tailor the uncertainty management to specific navigation tasks or spatial regions.
Risk-Aware Planning and Navigation:
@@
Risk-aware planning introduces measures to balance performance and safety by considering the potential risks associated with different actions (Li et al., 2022; Hakobyan et al., 2019). These techniques (Ono et al., 2015; Zhang et al., 2016) enable robots to make decisions that account for uncertainty in the environment (Hudack, 2016). In the context of RL, risk-sensitive approaches adjust the policy to avoid high-risk actions, often through modified reward functions or policy constraints (Chow et al., 2018; Lütjens et al., 2019; Srinivasan et al., 2020). Although effective in managing risk, these methods typically apply uniform risk thresholds and do not adapt to the task-specific uncertainty requirements that may vary across different navigation scenarios.
Task-Specific Uncertainty Management in Navigation:
@@
Integrating high-level task specifications into uncertainty management for navigation remains an open challenge. Some recent works propose large pre-trained models to interpret complex instructions and generate corresponding navigation behaviors (Shah et al., 2023b; Padalkar et al., 2023; Shah et al., 2023d; Huang et al., 2023; Shah et al., 2023c, a; Chang et al., 2023). While these approaches enable robots to understand and execute a wider range of tasks, they often lack a systematic method for representing and integrating task-specific uncertainty requirements into the navigation policy. This limitation reduces their effectiveness in complex environments where the importance of uncertainty can vary significantly across different regions.
In contrast to these prior works, our proposed GUIDE framework systematically and explicitly incorporates task-specific uncertainty requirements into both the planning and learning processes for navigation. By addressing the limitations of existing methods by treating uncertainty in a task-specific manner, our approach enables robots to dynamically adjust their navigation policies according to the varying importance of uncertainty.
3. The GUIDE Framework
Consider a robot operating in a continuous state space $S$, where each state $s\in S$ represents its pose. It can execute actions $a\in A$ from a continuous action space $A$. Assigned a navigation task $\tau$ specified by a natural language description, the robot must achieve objectives and adhere to constraints related to features of the environment. Our goal is to develop a navigation policy $\pi(a|s)$ that efficiently accomplishes the task while managing the robot’s state estimation uncertainty $u(s)$ in a task-aware manner. We aim to address the following problems using GUIDE:
@@ 3. The GUIDE Framework
    Interpret the task $\tau$ to determine acceptable levels of uncertainty across different regions of the environment, reflecting the relative importance of uncertainty reduction as dictated by the task requirements.
@@
    Synthesize a navigation policy that integrates this task-specific uncertainty information into the decision-making process. The policy should guide the robot to manage its uncertainty appropriately, prioritizing uncertainty reduction where necessary to ensure successful task execution, while efficiently pursuing task objectives in other areas.
Figure 2 presents an overview of our approach. The following sections detail the main components of GUIDE along with their implementation details.
3.1. Task-Specific Uncertainty Map (TSUM)

### v3

# GUIDEd Agents: Enhancing Navigation Policies through Task-Specific Uncertainty Abstraction in Localization-Limited Environments

@@
Abstract
Autonomous vehicles performing navigation tasks in complex environments face significant challenges due to uncertainty in state estimation. In many scenarios, such as stealth operations or resource-constrained settings, accessing high-precision localization comes at a significant cost, forcing the robots to rely primarily on less precise state estimates. Our key observation is that different tasks require varying levels of precision in different regions: a robot navigating a crowded space might need precise localization near obstacles but can operate effectively with less precise state estimates in open areas. In this paper, we present a planning method for integrating task-specific uncertainty requirements directly into navigation policies. We introduce Task-Specific Uncertainty Maps (TSUMs), which abstract the acceptable levels of state estimation uncertainty across different regions. TSUMs align task requirements and environmental features using a shared representation space, generated via a domain-adapted encoder. Using TSUMs, we propose Generalized Uncertainty Integration for Decision-Making and Execution (GUIDE), a policy conditioning framework that incorporates these uncertainty requirements into the robot’s decision-making process. We find that TSUMs provide an effective way to abstract task-specific uncertainty requirements, and conditioning policies on TSUMs enables the robot to reason about the context-dependent value of certainty and adapt its behavior accordingly. We show how integrating GUIDE into reinforcement learning frameworks allows the agent to learn navigation policies that effectively balance task completion and uncertainty management without the need for explicit reward engineering. We evaluate GUIDE on a variety of real-world robotic navigation tasks and find that it demonstrates significant improvement in task completion rates compared to baseline methods that do not explicitly consider task-specific uncertainty. Evaluation videos and code are available at https://guided-agents.github.io [^1] Marine navigation dataset will be made available on the website after the anonymous review period..
I Introduction
In complex environments where robots must balance task completion with resource usage, managing uncertainty in state estimation becomes a critical challenge.
@@ I Introduction
Figure 1: Illustration of our method applied to a robotic navigation task in a localization-limited environment. Top Left: The ASV is assigned a navigation task. Top Right: GUIDE interprets the task and generates a representation highlighting areas where the ASV needs higher positional certainty (dark blue). Bottom: The policy executed by the ASV. The white line represents the trajectory taken by the ASV towards the dock (green rectangle), the red dots indicate locations where the ASV actively reduces its state estimation uncertainty to satisfy task-specific requirements.
Consider an autonomous surface vehicle (ASV) conducting a mapping mission in stealth-critical settings where each GPS fix risks detection, or a ground robot operating in environments where high-precision localization requires costly computation. In such resource-constrained settings, the robot must not only complete its assigned task but also carefully manage when and where to use expensive localization operations. Importantly, constantly striving to reduce uncertainty through high-precision localization is neither necessary nor efficient for successful task execution. When operating in regions critical to the task objectives or near obstacles, precise localization becomes crucial for mission success. However, in areas less relevant to the immediate task goals, the robot can often tolerate higher positional uncertainty, allowing it to focus on efficient task execution rather than achieving pinpoint accuracy. This scenario underscores a fundamental challenge in robotic navigation: in environments with limited access to high-precision localization, the robot must balance task completion with uncertainty management, where the acceptable level of uncertainty at any location is inherently task-specific and context-dependent.
Traditional approaches to navigation in environments with limited access to high-precision localization often fall into two categories: minimizing uncertainty universally through frequent use of expensive localization [47, 12, 37, 13] or enforcing fixed uncertainty thresholds across the environment [7, 19, 52]. While these strategies may be effective in settings where localization resources are abundant, they become impractical when high-precision state estimation comes at a cost. For instance, in stealth-critical missions, universally enforcing precise localization through frequent GPS usage would compromise the robot’s covertness, even in regions where such precision isn’t necessary for task execution.
@@
Existing methods also struggle to balance resource usage with task requirements. While some approaches attempt to reduce localization costs by considering task performance [65, 63], they typically apply a uniform trade-off across all situations. This fails to account for how the value of precise localization varies across different phases of the mission. Furthermore, these methods often require extensive reward engineering to capture the complex relationship between task objectives and localization costs [77, 5].
The fundamental issue lies in the disconnect between task requirements and resource-constrained uncertainty management. Current frameworks treat localization decisions independently from task specifications, leading to either excessive resource consumption or compromised task performance when different regions demand different levels of certainty.
@@
Statement of contributions: With the aim of bridging the gap between high-level task descriptions and low-level uncertainty handling, we propose the concept of a Task-Specific Uncertainty Map (TSUM). A TSUM serves as an intermediate abstraction that represents the acceptable levels of state estimation uncertainty across different regions of the environment for a specific task. We present a CLIP-based approach for computing TSUMs that effectively aligns visual environmental features with textual task specifications in a shared embedding space.
@@
Building upon TSUM, we introduce Generalized Uncertainty Integration for Decision-Making and Execution (GUIDE), a policy conditioning framework that incorporates TSUMs into navigation policies. The core proposition of GUIDE is that by conditioning navigation policies on these task-specific uncertainty requirements, robots can better reason about the context-dependent value of certainty. This enables the generation of policies that dynamically adapt their behavior based on the varying uncertainty requirements across different regions of the environment.
@@
We evaluate GUIDE in the context of marine autonomy using Autonomous Surface Vehicles (ASVs). We chose marine navigation as our experimental domain for three key reasons: (i) marine navigation presents inherently greater challenges than ground navigation due to environmental uncertainty from waves, currents, and weather conditions, (ii) it enables evaluation through real-world, ‘in-the-wild’ deployment rather than controlled laboratory settings, and (iii) it represents a less-explored area in robotics with limited datasets and established benchmarks. Our experimental results show significant improvements in task performance when compared to state-of-the-art methods that do not explicitly consider task-specific uncertainty requirements.
II Related Works
Uncertainty Modeling in Robotic Navigation
@@ II Related Works
Uncertainty management is a fundamental aspect of robotic navigation, where robots must make decisions based on imperfect information about their state and the environment [47, 57]. Probabilistic techniques such as Bayesian filters [15, 25, 39] and particle filters [16, 66, 11] have been widely used to enable robust localization and mapping in stochastic environments. These methods allow robots to estimate their position and navigate effectively despite sensor noise and environmental uncertainties. However, traditional approaches often treat uncertainty uniformly across the environment, without considering how different regions may require varying levels of certainty depending on the navigation task. While these methods effectively estimate and track uncertainty, they are not designed for scenarios where reducing uncertainty comes at a significant cost. In resource-constrained environments where high-precision localization is limited, simply estimating uncertainty is insufficient - robots must also reason about when to invest in reducing it.
Reinforcement Learning for Navigation
@@
Reinforcement Learning (RL) has been successfully applied to robotic navigation tasks, enabling agents to learn navigation policies through interaction with the environment [69, 70, 75]. Standard RL algorithms have been used to train robots to navigate in various settings, from simulated environments to real-world scenarios [31, 74, 67]. While these methods can learn effective policies in controlled environments [79, 46, 27], they often perform suboptimally in real-world scenarios due to factors like partial observability, dynamic changes, and varying levels of uncertainty [8, 38]. Additionally, standard RL typically does not account for uncertainty explicitly, which can hinder performance in complex environments where uncertainty plays a significant role. Moreover, lot of these methods typically assume unbounded access to state information, making them poorly suited for environments where precise state estimation is a limited resource that must be strategically utilized.
@@
Figure 2: Overview of GUIDE. The framework consists of two phases: Embedding Alignment (Pre-training) and Deployment. In the embedding alignment phase, semantic and spatial embeddings are fine-tuned and aligned using a CLIP-based model with contrastive learning. During deployment, TSUMs are generated from natural language task descriptions and environmental data, and navigation policies are conditioned on these TSUMs to manage uncertainty in a task-aware manner.
Task-Specific Navigation Policies
@@
Recent advances have focused on developing navigation policies that are conditioned on specific tasks or goals [51, 10, 41, 36]. Task-conditioned RL allows agents to adapt their behavior based on high-level task specifications, improving flexibility and generalization. In navigation, this has been explored by training policies capable of handling objectives, adjusting the robot’s behavior according to the task [78, 68, 40]. These approaches often neglect the varying importance of uncertainty management [9, 26] across different tasks and environments.
Uncertainty-Aware Reinforcement Learning in Navigation
@@
Incorporating uncertainty into RL for navigation has been investigated to enhance exploration, robustness, and safety [62, 75]. Some approaches introduce uncertainty penalization directly into the reward function, encouraging agents to avoid high-uncertainty regions [14, 76, 33]. Others utilize Bayesian RL methods to model uncertainty in value estimation, improving decision-making under uncertainty [3, 1]. Bootstrapped ensembles maintain multiple value function estimates to capture epistemic uncertainty, leading to more informed exploration strategies [43, 2, 28, 71]. While these methods consider uncertainty, they often do so globally or uniformly [48, 30, 44] across the environment and do not tailor the uncertainty management to specific navigation tasks or spatial regions. This uniform treatment of uncertainty becomes particularly problematic in resource-constrained settings, where robots have limited opportunities to reduce uncertainty. These methods lack mechanisms to reason about when consuming these limited resources would provide the most value for task completion.
Risk-Aware Planning and Navigation
@@
Risk-aware planning introduces measures to balance performance and safety by considering the potential risks associated with different actions [32, 18]. These techniques [42, 73] enable robots to make decisions that account for uncertainty in the environment [22]. In the context of RL, risk-sensitive approaches adjust the policy to avoid high-risk actions, often through modified reward functions or policy constraints [6, 35, 64]. Although effective in managing risk, these methods typically apply uniform risk thresholds and do not adapt to the task-specific uncertainty requirements that may vary across different navigation scenarios.
Task-Specific Uncertainty Management in Navigation
@@
Integrating high-level task specifications into uncertainty management for navigation remains an open challenge. Some recent works propose large pre-trained models to interpret complex instructions and generate corresponding navigation behaviors [59, 45, 61, 21, 60, 58, 4]. While these approaches enable robots to understand and execute a wider range of tasks, they often lack a systematic method for representing and integrating task-specific uncertainty requirements into the navigation policy. This limitation reduces their effectiveness in complex environments where the importance of uncertainty can vary significantly across different regions.
In contrast to prior works, we place our research at the intersection of task-specific navigation and resource-constrained uncertainty management. Our proposed GUIDE framework addresses a critical gap in existing literature: how to effectively navigate in environments where high-precision localization is a limited resource that must be strategically utilized. While previous approaches either assume unlimited access to precise state estimation or treat uncertainty uniformly across tasks, GUIDE systematically incorporates both task requirements and localization constraints into the navigation policy. By providing a principled way to reason about when and where to use costly localization operations, our approach enables robots to dynamically balance task completion with resource usage, adapting their behavior based on the varying importance of uncertainty across different regions of the environment.
III The GUIDE Framework
Consider a robot operating in a continuous state space $\mathcal{S}$, where each state $s\in\mathcal{S}$ encodes the robot’s pose. The robot can execute control inputs from a continuous action space $\mathcal{A}$. Operating in a localization-limited environment, the robot’s state estimate is subject to uncertainty due to sensor noise, communication constraints, and environmental factors. While the robot can access high-precision localization readings, doing so incurs significant costs—for instance, an autonomous surface vehicle (ASV) conducting stealth-critical missions where each GPS fix risks detection, or a ground robot in environments where precise localization demands substantial computational resources.
We assume the availability of a visual representation of the operating environment a priori, such as an overhead map or imagery covering the region of interest. This assumption is justified in many real-world applications where survey data or geographic information systems provide baseline knowledge of the environment (e.g., topography, known obstacles).
A navigation task $\tau$, specified in natural language, describes the robot’s objectives and associated constraints with respect to environmental features. For example, $\tau$ might require the robot to “navigate to the dock while avoiding right quadrant”. The essential characteristic is that these tasks reference spatial or semantic features describing the environment.
Objective: Synthesize a navigation policy $\pi(a\mid s)$ that enables task completion while explicitly reasoning about the robot’s state-estimation uncertainty $u(s)$. The policy must balance two competing demands: efficient task execution and strategic uncertainty management. This balance is particularly crucial in resource-constrained settings where reducing uncertainty comes at a significant cost.
To achieve this objective, the GUIDE framework addresses two fundamental challenges in uncertainty-aware navigation:
@@ III The GUIDE Framework
    Task Interpretation: Given a natural language task $\tau$, determine acceptable levels of uncertainty across different regions of the environment. This interpretation captures how important precise state estimation is for each area of the operating space, based on task requirements and environmental features.
@@
    Policy Synthesis: Generate a navigation policy that incorporates task-specific uncertainty requirements into its decision-making process. This policy should prioritize uncertainty reduction in critical regions while efficiently pursuing task objectives where precise localization is less crucial.
Fig. 2 presents an overview of the GUIDE framework. In the following sections, we detail how these components are implemented within a unified navigation architecture.
III-A Task-Specific Uncertainty Map (TSUM)